288 286 295 267 332 ?
Answer by Edwin McCravy(20056)

288 286 295 267 332 ?
Begin with 288
Subtract 13+1, which equals 1+1 = 2
Get 288-2 = 286

Take 286
Add 23+1, which equals 8+1 = 9
Get 286+9 = 295

Take 295
Subtract 33+1, which equals 27+1 = 28
Get 295-28 = 267

Take 267
Add 43+1, which equals 64+1 = 65
Get 267+85 =  332 

Take 332
Subtract 53+1, which equals 125+1 = 126
Get 332-126 = 206

Edwin